OpenStudy (anonymous):

integral from ln5 to ln10

OpenStudy (anonymous):

that isn't the whole question let me finish it

OpenStudy (anonymous):

\[ \int\limits_{\ln5}^{\ln10}\pi(e ^{x/2})^{2}-\pi(e ^{-x})^{2} dx\]